# NIMO: a Nonlinear Interpretable MOdel

### v1

@@ Abstract
Neural networks (NNs) have achieved tremendous success over the past decade, yet they are still extremely difficult to interpret. In contrast, linear models are less expressive but offer inherent interpretability. Linear coefficients are interpretable as the marginal effect of a feature on the prediction, assuming all other features are kept fixed. To combine the benefits of both approaches, we introduce NIMO (Nonlinear Interpretable MOdel). The key idea is to define a model where the NN is designed to learn non-linear corrections to the linear model predictions, while also maintaining the original interpretability of the linear coefficients. Relevantly, we develop an optimization algorithm based on profile likelihood that elegantly allows for optimizing over the NN parameters while updating the linear coefficients analytically. By relying on adaptive ridge regression we can easily incorporate sparsity constraints as well. We show empirically that we can recover the underlying linear coefficients while significantly improving the predictive accuracy. Compared to other hybrid interpretable approaches, our model is the only one that actually maintains the same interpretability of linear coefficients as in linear models. We also achieve higher performance on various regression and classification settings.
@@ 1 Introduction
Over the course of the past decade machine learning research has witnessed enormous success thanks to neural networks. Applications range from computer vision, natural language processing, speech recognition, robotics, finance and many more. Along with the development of these remarkable capabilities came the desire for understanding the underlying decision-making process. However, neural networks still lack a clear interpretability, which prevents their widespread use in high-stake areas like healthcare. Most research has focused on how to address this problem, either by developing intrinsically interpretable models or by providing post-hoc explanations of model predictions.
In this work, we focus on the design of an intrinsically interpretable model by employing the inherent transparency of linear models. Because of the linear relationship between the input data and the output predictions, linear models are simple and their decision-making process is transparent. The sign of a linear coefficient indicates the direction of the relationship (positive or negative), and its magnitude reflects the significance of that feature on the prediction. However, the simplicity of linear models constrains their expressiveness. In contrast, neural networks have shown great ability in learning complex feature maps from the input data in order to predict the output. While this lead to an enormous success in many applications, NNs lack interpretability of the predictions in terms of the input features, hence the term “black-box” models. This is particularly evident in the medical domain where, in order to make informed decisions, practitioners need to understand which factors lead to a particular model outcome.
To solve this dilemma, we introduce the Nonlinear Interpretable MOdel (NIMO), which inherits the expressivity of neural networks and also maintains the transparency and interpretability of linear models. More specifically, we start from a linear model and from the associated linear coefficients $\bm{\beta}$ that provide a clear interpretability at a population level. Then, we adjust each $\beta_{j}$ by a multiplicative factor $h_{j}$ according to the individual point $\bm{x}$. We learn such correction terms with a neural network $\text{NN}_{\bm{u}}$ so that the model can learn nonlinear interaction terms beyond the linear ones:
$y=\sum_{j=1}^{p}x_{j}\beta_{j}h_{j}\quad\text{with}\quad h_{j}=\text{NN}_{\bm{u}}(\bm{x})\;.$ (1)
If the neural network has a suitable design, the resulting model maintains the interpretability of linear coefficients $\bm{\beta}$ at a global level while also providing a nonlinear correction at the local level, depending on the data point $\bm{x}$. In Figure 1 we illustrate how the proposed approach shares the same high-level structure of a linear model while also integrating nonlinear corrections through neural networks. In theory we would need one neural network per feature, but in practice we can use a single network with the feature appended with positional encoding. In Section 3 we show that, with simple requirements on the neural network, we can preserve the original interpretability of the linear coefficients $\bm{\beta}$. Since the network acts as a nonlinear correction to the linear coefficients $\bm{\beta}$, the joint optimization of $\bm{\beta}$ and $\bm{u}$ might not be trivial in practice. To address this issue, we develop an optimization method based on the profile likelihood which allows for sparsity constraints as well. Specifically, we alternate a gradient step over the neural network parameters with an analytical update over the linear coefficients $\bm{\beta}$. Lastly, we show that we can extend the present framework to generalized linear models.
As commonly done in the literature, we further enforce interpretability by means of sparsity. When fewer components (features or connections) are active, it is easier to understand how the information leading to some prediction is processed. In the linear part of our model, we employ L1 regularization to enforce the sparsity in the coefficients $\bm{\beta}$. In the nonlinear part originating from the network, we resort to the group L2 regularization on the weight matrix of the first fully connected (fc) layer. We also employ a novel noise injection technique to the output of the first fc layer in order to encourage the selection of the most relevant features.
@@
  • We introduce a nonlinear interpretable model (NIMO) that combines the expressive power of neural networks with the well-established interpretability of linear models.
  • We introduce an optimization algorithm that allows to jointly optimize the linear coefficients and the nonlinear contributions of the neural network.
  • On synthetic data, we show that we retrieve the underlying linear coefficients, learn the correct nonlinearities and outperform state-of-the-art hybrid models. On real data, we demonstrate NIMO’s interpretability by showing that we recover sparse coefficients while achieving performance on par with neural network approaches.
@@ 2 Related Work
Interpretability in machine learning. Interpretability refers to the capacity to express what a model has learned and the factors influencing its outputs in a manner that is clear and understandable to humans. In the literature, the two terms interpretability and explainability are sometimes used interchangeably. However, they have subtle yet important differences. Explainability focuses on providing reasons for the model’s output. Most of the methods, especially in deep learning, provide post-hoc explanations. The most popular ones include LIME (Ribeiro et al., 2016), SHAP (Lundberg and Lee, 2017) and Grad-CAM (Selvaraju et al., 2020). On the other hand, interpretability focuses on understanding the inner mechanisms and decision processes of the model, and most of the methods involve inherently understandable models, such as decision trees (Wu et al., 2017, 2021), Lasso (Tibshirani, 1996) and Explainable Boosting Machines (Lou et al., 2013). Often, simple interpretability comes at the cost of reduced expressivity. The literature on interpretable machine learning is vast and we refer readers to (Rudin et al., 2022; Molnar, 2025) for a more in-depth review. In this work, we develop a model that has a straightforward interpretability and that, at the same time, is very expressive and flexible.
Interpretability of linear models. Linear models are inherently interpretable. They model the relationship between features and the target variable as a linear combination, making the effect of each feature easy to understand and quantify (Molnar, 2025). For example, consider a linear regression model used to predict the risk of heart disease based on a set of clinical features. Then, the coefficient associated with a specific clinical feature represents the change in the predicted risk when that feature increases by one unit, assuming all other feature values remain fixed. Methods like Lasso regularization (Tibshirani, 1996) can be further applied to achieve sparse feature selection and focus on the most relevant features. However, linear models have limited expressiveness and fail to capture complex relationships among the features. In this work, we maintain the interpretability of linear models and, at the same time, we increase the model expressiveness through a neural network.
Hybrid models. Several works have proposed to combine linear models with neural networks to achieve more interpretable models. However, none of them actually retain the original interpretability of linear coefficients as in linear models. LassoNet (Lemhadri et al., 2021) is the first approach to integrate a neural network with Lasso regression. It allows a feature to be used in the neural network only if its corresponding Lasso coefficient is nonzero. This constraint ensures that both the neural network’s nonlinear component and the Lasso linear component operate on the same subset of features. Even though this allows for feature selection, deactivating the features that are deemed unimportant in the linear part significantly constrains the capacity of the neural network. Contextual Lasso (Thompson et al., 2023) is a type of varying coefficient model (Hastie and Tibshirani, 1993). The linear regression coefficients are context-dependent and generated by a neural network. This structure provides flexibility but comes at the cost of sacrificing global interpretability of the coefficients, which is key for the interpretability of linear models. In this work, we develop an inherently interpretable model called NIMO, which maintains the global interpretability of linear models while also learning nonlinear corrections to the linear predictions.
@@ 3 Proposed approach
In Section 3.1 we introduce our non-linear interpretable model, namely NIMO, and argue why we can still retain the full interpretability of linear models while also allowing for non-linearities. Then, in Section 3.2 we propose an algorithm to elegantly train the model on the neural network parameters while analytically updating the linear coefficients. In Section 3.3 we show how to find relevant nonlinear interactions by means of sparsity. Lastly, In Section 3.4 we discuss the limitations.
@@ 3.1 A Nonlinear Interpretable Model
Let $X\in\mathbb{R}^{n\times d}$ be $n$ $d$-dimensional observations and let $\bm{y}\in\mathbb{R}^{n}$ be the targets. In the following, we assume the data $X$ to be standardized, i.e. that each feature has zero mean and unit standard deviation. A linear model is simply defined by the linear coefficients $\bm{\beta}\in\mathbb{R}^{d}$ and a bias term $\beta_{0}\in\mathbb{R}$ as $y_{i}\approx\beta_{0}+\sum_{j=1}^{d}x_{j}\beta_{j}$. In linear models the coefficient $\beta_{j}$ is a proxy for the importance of the $j$-th feature. More precisely, the linear coefficient $\beta_{j}$ can be interpreted as the additive effect of the $j$-th feature on the output with all other variables at baseline level. The idea behind our approach is to maintain the interpretability of global linear coefficients while allowing for individual nonlinear corrections to the linear predictions. In order to accomplish this, we define our approach starting from a linear model and by multiplying the linear coefficients $\beta_{j}$ by a nonlinear correction term $h_{j}$ that depends on the individual data point $\bm{x}$. The model is defined as follows:
$f(\bm{x})=\beta_{0}+\sum_{j=1}^{d}x_{j}\beta_{j}\underbrace{(1+g_{\bm{u}_{j}}(\bm{\bm{x}_{-j}}))}_{h_{j}}\quad\text{s.t.}\quad g_{\bm{u}_{j}}(\mathbf{0})=0\;,$ (2)
where $\bm{\bm{x}_{-j}}$ is the vector $\bm{x}$ without the $j$-th component and where $g_{\bm{u}_{j}}:\mathbb{R}^{d-1}\mapsto\mathbb{R}$ are $d$ different neural networks, each parametrized by $\bm{u}_{j}$. We provide an illustration of the proposed approach in Figure 2. To make sure that our model in Eq. (2) fulfills the intended interpretability we took care of three aspects. First, the network $g_{\bm{u}_{j}}$ needs to act as a non-linear correction to the linear prediction. This is easily achieved by designing the neural network as an additive contribution to the linear prediction. We multiply the coefficient $\beta_{j}$ by $h_{j}=1+g_{\bm{u}_{j}}$, which renders the additive effect: $f(\bm{x})=\beta_{0}+\sum_{j=1}^{d}x_{j}\beta_{j}+\sum_{j=1}^{d}x_{j}\beta_{j}g_{\bm{u}_{j}}(\bm{\bm{x}_{-j}})$. Second, the neural network $g_{\bm{u}_{j}}$ cannot depend on the $j$-th feature itself. Otherwise, $x_{j}$ would contribute to the prediction not only through the linear term $\beta_{j}$ but also through the network itself, which would change the behavior and interpretation of the coefficient $\beta_{j}$. Third, we make sure that the non-linear contribution coming from the neural network keeps the feature at baseline level (as in the linear part), i.e. $g_{\bm{u}_{j}}(\bm{\bm{x}_{-j}}=\bm{0})=\bm{0}$. It is easy to see that when $\bm{\bm{x}_{-j}}=\bm{0}$ the model reduces to classical linear regression.
@@
Figure 2: NIMO architecture. The model consists of global linear coefficients $\beta_{j}$ modulated by the individual data point $\bm{x}$ through $h_{j}$. In particular, $h_{j}$ depends on $\bm{\bm{x}_{-j}}$, i.e. all features except the $j$-th.
Model interpretability
At a high level, the proposed approach resembles a linear model where each coefficient is given by the product $\beta_{j}h_{j}$. Intuitively, $\beta_{j}$ is independent of the data point $\bm{x}$ and acts at a population level. In contrast, $h_{j}$ provides a nonlinear correction to $\beta_{j}$ and depends on the individual data point $\bm{x}$ (more precisely $\bm{\bm{x}_{-j}}$) through $g_{\bm{u}}$. This construction allows to keep the global interpretability as in a linear model while also allowing for a local nonlinear correction. To see why this is relevant, consider a healthcare application where we want to predict a treatment outcome based on some clinical data. The resulting linear coefficients represents at a population level the contribution of each feature to the prediction. However, the global explanation might not be a good fit for all patients, which is a limitation of linear models. For some patients the outcome might be better explained by a different coefficient, which we can learn by multiplying the global $\beta_{j}$ by a local $h_{j}$ specific to that patient. Overall, the proposed approach provides a population-wise interpretation of linear coefficients while still allowing for interpretable nonlinear and local corrections.
@@
The proposed model in Eq. (2) requires $d$ different neural networks, which can require significant memory in high-dimensional settings. Instead, we use a single shared network $g_{\bm{u}}$ and employ positional encoding to make the network aware of which component is being queried. This can be easily done by appending the positional encoding of the $j$-th component to the input $\bm{\bm{x}_{-j}}$. In practice, $g_{\bm{u}}$ takes the full vector $\bm{x}$ as input and we can simply mask out the $j$-th component of $\bm{x}$ to get $g_{\bm{u}}(\bm{\bm{x}_{-j}})$. Furthermore, we enforce $g_{\bm{u}}(\bm{\bm{x}_{-j}}=\bm{0})=0$ by design by simply subtracting the mean prediction at each forward pass: $g_{\bm{u}}(\bm{\bm{x}_{-j}})\coloneqq g_{\bm{u}}(\bm{\bm{x}_{-j}})-g_{\bm{u}}(\bm{0})$.
3.2 Training through profile likelihood with adaptive ridge regression
For illustrative purposes, in this section we provide a high-level description of the optimization algorithm used to train NIMO. We initially consider a regression setting and later extend the same reasoning to generalized linear models, in particular for the logistic regression case.

### v2

@@ Abstract
Deep learning has achieved remarkable success across many domains, but it has also created a growing demand for interpretability in model predictions. Although many explainable machine learning methods have been proposed, post-hoc explanations lack guaranteed fidelity and are sensitive to hyperparameter choices, highlighting the appeal of inherently interpretable models. For example, linear regression provides clear feature effects through its coefficients. However, such models are often outperformed by more complex neural networks (NNs) that usually lack inherent interpretability. To address this dilemma, we introduce NIMO, a framework that combines inherent interpretability with the expressive power of neural networks. Building on the simple linear regression, NIMO is able to provide flexible and intelligible feature effects. Relevantly, we develop an optimization method based on parameter elimination, that allows for optimizing the NN parameters and linear coefficients effectively and efficiently. By relying on adaptive ridge regression we can easily incorporate sparsity as well. We show empirically that our model can provide faithful and intelligible feature effects while maintaining good predictive performance.
@@ 1 Introduction
Over the past decade, neural networks have achieved remarkable success across domains, from computer vision (Krizhevsky et al., 2012) to natural language processing (Vaswani et al., 2017). At the same time, their deployment in high-stakes domains such as healthcare has created a pressing demand for interpretability (Esteva et al., 2017). However, neural networks in general lack interpretability of the predictions in terms of the input features, hence the term “black-box” models. In contrast, classical models such as linear regression and decision trees are often considered as highly interpretable because of their transparent structure and simple decision rules, but they can lack predictive power on complex, high-dimensional, or highly nonlinear tasks (Hastie, 2009; Breiman et al., 2017). This tension between accuracy and interpretability has motivated much recent work on models and methods that combine the expressive power of neural networks with the transparency of simpler models (Lemhadri et al., 2021; Thompson et al., 2023; Wu et al., 2017; 2021). Meanwhile, many model-agnostic, post-hoc explainers have been proposed, such as SHAP (Lundberg & Lee, 2017) and LIME (Ribeiro et al., 2016), providing instance-level feature attributions for arbitrary predictors. But these post-hoc explanations are approximations and may depend on the choices of parameters, such as the background distribution and kernel width. Among various ways to achieve interpretability, feature effects (Scholbeck et al., 2019) provide a direct, quantitative description of how each feature influences the prediction, something post-hoc methods can approximate but do not guarantee.
We use the term feature effects in a broad sense, referring to how changes in input features influence model predictions. A common formalization of this idea is through marginal effects (Nguyen, 2020), which define feature effects as the direction and magnitude of the change in prediction due to an infinitesimal change in a feature value. Linear models offer immediately intelligible marginal effects through their coefficients (Molnar, 2025), but are outperformed by more complex nonlinear models, particularly neural networks. Unfortunately, these more expressive models typically lose the property of providing such inherently intelligible effects. In such models, marginal effects must instead be obtained by computing derivatives of the fitted function with respect to the inputs, usually through computationally expensive automatic differentiation (autograd) frameworks (Paszke et al., 2017). To bridge this gap, we propose an intrinsically interpretable hybrid model that preserves intelligible marginal effects while retaining the expressivity of neural networks. Specifically, we start from a linear model $y=\sum_{j}x_{j}\beta_{j}$, and adjust each $\beta_{j}$ by a multiplicative factor $h_{j}$, which is the output of a neural network $\text{NN}_{\bm{u}_{j}}(\mathbf{x})$ evaluated at the input instance $\mathbf{x}$:
$y=\sum_{j=1}^{p}x_{j}\beta_{j}h_{j}(\mathbf{x})\quad\text{with}\quad h_{j}(\mathbf{x})=\text{NN}_{\bm{u}_{j}}(\mathbf{x})\;.$ (1)
In Figure 1, we illustrate how the proposed approach retains the high-level structure of a linear model while integrating per-instance nonlinear corrections through neural networks.
Within this framework, we distinguish between local explanations and global interpretations. Local explanations describe how changes in a feature influence the prediction for a specific instance, enabled here by the per-instance corrections $h_{j}(\mathbf{x})$ from the neural network. Global interpretations, in contrast, summarize the overall influence of a feature across the dataset, which we later formalize through the marginal effects at the mean (MEM). In our model, these coincide exactly with the global linear coefficients. To illustrate, consider a medical application where we aim to estimate a patient’s risk of developing a disease. A local explanation would ask: given this patient’s age, weight, and other characteristics, how does a small change in age affect their individual risk? A global interpretation instead asks: if all other variables are held fixed, how does age in general influence disease risk across the population? For linear models, these two perspectives coincide, but in our framework they differ due to potential nonlinear feature interactions. Our network is designed so that the model preserves the classical interpretation of the linear coefficients $\bm{\beta}$, while simultaneously providing per-instance corrections through $h_{j}(\mathbf{x})$. This enables us to unify global summaries via MEM with local explanations at the instance level within a single coherent framework.
Since the network parameters $\bm{u}$ and the linear coefficients $\bm{\beta}$ are tightly coupled, jointly optimizing them is non-trivial. To address this, we develop an optimization method based on parameter elimination. Specifically, we derive a closed-form expression for the linear coefficients $\bm{\beta}$ as a function of the NN parameters $\bm{u}$. Substituting this expression back into the objective reduces the problem to optimizing only over the NN parameters. To further enhance interpretability, we impose sparsity by formulating the problem as adaptive ridge regression (Grandvalet, 1998). In the network, we further apply group $\ell_{2}$ regularization (Yuan & Lin, 2006) to the weight matrix of the first fully connected layer, encouraging feature-level sparsity and clarifying how inputs contribute to predictions.
@@
  • We introduce a nonlinear interpretable model that combines the expressivity of neural networks with the interpretability of linear models, providing intelligible global interpretability via marginal effects at the mean (MEM) and per-instance corrections through the network.
  • We introduce an optimization algorithm via parameter elimination that effectively and efficiently optimizes the linear coefficients and the neural network parameters .
  • We systematically evaluate NIMO on synthetic and real datasets, verifying that it recovers correct feature effects, achieves accurate per-instance explanations, and offers a favorable trade-off between interpretability and predictive performance compared to existing methods.
@@ 2 Related Work
Interpretability in machine learning. Interpretability refers to the capacity to express what a model has learned and the factors influencing its outputs in a manner that is clear and understandable to humans. In the literature, the two terms interpretability and explainability are sometimes used interchangeably. However, they have subtle yet important differences. Explainability focuses on providing reasons for the model’s output. Most of the methods, especially in deep learning, provide post-hoc explanations. The most popular ones include LIME (Ribeiro et al., 2016), SHAP (Lundberg & Lee, 2017) and Grad-CAM (Selvaraju et al., 2020). On the other hand, interpretability focuses on understanding the inner mechanisms and decision processes of the model, and most of the methods that claimed to be interpretable involve inherently understandable models, such as decision trees (Wu et al., 2017; 2021), Lasso (Tibshirani, 1996) and Explainable Boosting Machines (Lou et al., 2013). The literature on interpretable machine learning is vast and we refer readers to Rudin et al. (2022) and Molnar (2025) for a more in-depth review.
Linear Models and Input Feature Effects. Coefficients from linear regression models provide inherently interpretable feature effects, directly quantifying how changes in input features influence predictions (Nguyen, 2020; Molnar, 2025). This transparency makes linear models especially attractive in domains where interpretability is critical. Regularization methods such as the Lasso (Tibshirani, 1996) further enhance interpretability by selecting the most relevant input features. However, linear models cannot capture nonlinearities or interactions without complicated feature engineering, e.g., basis expansion, which limits their ability to reflect complex data-generating processes. This motivates our approach, which preserves the interpretability of linear coefficients while extending their expressiveness to capture more complex data patterns.
Hybrid models. Several works have proposed combining linear models with neural networks to improve interpretability. Neural Additive Models (NAMs) (Agarwal et al., 2021) extend generalized additive models by learning one neural network per feature, whose outputs are summed to form the prediction. This design preserves interpretability at the feature level but does not have the ability to capture feature interactions. LassoNet (Lemhadri et al., 2021) integrates a neural network with Lasso regression by enforcing that a feature can be used in the nonlinear part only if its corresponding linear coefficient is nonzero. Interpretable Mesomorphic Networks (IMN) (Kadra et al., 2024) use a hypernetwork to predict linear coefficients on a per-instance basis. While this increases flexibility, it sacrifices global interpretability of the coefficients, making it difficult to recover clear baseline effects. In contrast, our proposed model (NIMO) preserves the global interpretability of linear coefficients while augmenting them with nonlinear corrections. This unifies global summaries via MEM with instance-level explanations within a single coherent framework.
@@ 3.1 A Nonlinear Interpretable Model
Let $X\in\mathbb{R}^{n\times d}$ be $n$ observations with $d$ dimensions and let $\mathbf{y}\in\mathbb{R}^{n}$ be the targets. In the following, we assume the data $X$ to be standardized, i.e. that each feature has zero mean and unit standard deviation. A linear regression model is simply defined as $y=\beta_{0}+\sum_{j=1}^{d}x_{j}\beta_{j}$. In linear models the coefficient $\beta_{j}$ can be interpreted as the additive effect of the $j$-th feature on the output when all other feature values remain fixed. The idea of our model is to maintain the intelligible global interpretation at the population level while allowing for local explanations at the per-instance level. To do so, we define our approach starting from a linear model and multiplying the linear coefficients $\beta_{j}$ by a nonlinear correction term $h_{j}$ that depends on the data point $\mathbf{x}$:
$f(\mathbf{x})=\beta_{0}+\sum_{j=1}^{d}x_{j}\beta_{j}\underbrace{(1+g_{\bm{u}_{j}}(\mathbf{\mathbf{x}_{-j}}))}_{h_{j}}\quad\text{s.t.}\quad g_{\bm{u}_{j}}(\mathbf{0})=0\;,$ (2)
where $\mathbf{\mathbf{x}_{-j}}$ is the vector $\mathbf{x}$ without the $j$-th component and $g_{\bm{u}_{j}}(\cdot)$ are $d$ different scalar-valued functions defined by neural networks, each parametrized by $\bm{u}_{j}$. We provide an illustration of the proposed approach in Figure 2. Due to the specific design of our model, it exhibits the following properties. First, by explicitly excluding the $j$-th feature as input of the neural network, i.e. $g_{\bm{u}_{j}}(\mathbf{\mathbf{x}_{-j}})$, $x_{j}$ contributes to the prediction only through the linear term $\beta_{j}$, preserving the behavior and interpretation of the coefficient $\beta_{j}$. Second, as we assume the data is standardized, the mean value of the features is zero. Therefore, since by construction $g_{\bm{u}_{j}}(\mathbf{0})=0$, when all other features are fixed at the mean value (i.e. at zero), the model reduces to a linear model. In other words, the marginal effect of feature $j$ is exactly $\beta_{j}$, as in a purely linear model. This ensures that the nonlinear correction term does not alter the global interpretability of the linear coefficients.
@@
Figure 2: NIMO architecture. The model consists of global linear coefficients $\beta_{j}$ modulated by the individual data point $\mathbf{x}$ through $h_{j}$. In particular, $h_{j}$ depends on $\mathbf{\mathbf{x}_{-j}}$, i.e. all features except the $j$-th.
@@
A naive implementation of Eq. 2 requires $d$ separate neural networks, which is impractical in high-dimensional settings. To address this, we use a single shared network $g_{\bm{u}}$ and provide it with positional information to distinguish between features. Concretely, we append a positional encoding of the index $j$ to the masked input vector $\mathbf{\mathbf{x}_{-j}}$, where the $j$-th component of $\mathbf{x}$ is set to zero. This way, $g_{\bm{u}}$ can learn feature-specific corrections while reusing parameters across all features. Finally, to enforce the interpretability constraint $g_{\bm{u}}(\mathbf{0})=0$, we simply subtract the mean prediction in each forward pass: $g_{\bm{u}}(\mathbf{\mathbf{x}_{-j}})\coloneqq g_{\bm{u}}(\mathbf{\mathbf{x}_{-j}})-g_{\bm{u}}(\bm{0})$.
3.2 Interpretability via marginal effects
Following Nguyen (2020), the marginal effect (ME) of feature $j$ at input $\mathbf{x}$ is defined as
$\mathrm{ME}_{j}=\frac{\partial f(\mathbf{x})}{\partial x_{j}},$ (3)
for any model $f(\mathbf{x})$. By definition, $\mathrm{ME}_{j}$ measures the instantaneous change in prediction when varying $x_{j}$. However, marginal effects in nonlinear models are not constant, as they depend on the other input features. One way to summarize them is by computing marginal effects at the mean (MEM), which estimates marginal effects at the average values of the input features:
$\mathrm{MEM}_{j}=\left.\frac{\partial f(\mathbf{x})}{\partial x_{j}}\right|_{\mathbf{\mathbf{x}}=\bar{\mathbf{x}}}.$ (4)
Based on the definition, we can compute the marginal effects for NIMO as below:
$\mathrm{ME}_{j}$ $=\beta_{j}\big(1+g_{\bm{u}_{j}}(\mathbf{x}_{-j})\big)+\sum_{i\neq j}x_{i}\beta_{i}\frac{\partial g_{\bm{u}_{i}}(\mathbf{x}_{-i})}{\partial x_{j}}\;,$
$\mathrm{MEM}_{j}$ $=\beta_{j}\;.$
Note that the mean value of input features $\bar{\mathbf{x}}=\mathbf{0}$ after standardization. Due to the design of our model, $\mathrm{MEM}_{j}$ coincides with the linear coefficient. In principle, such marginal effects can also be computed via automatic differentiation for other differentiable models, but these do not necessarily yield interpretable results. In contrast, for NIMO, the MEMs directly correspond to the linear coefficients $\bm{\beta}$, capturing the additive population-level contribution of each feature.
3.3 Training via parameter elimination
For illustrative purposes, in this section we provide a high-level description of the optimization algorithm used to train NIMO. More detailed training procedure can be found in Appendix B.